A body is thrown with speed 20m/s vertically upward. It will return to thrower's hand after a time?(g=10m/s square)
1
Expert's answer
2012-05-22T08:01:11-0400
If it was thrown straight vertically upward it surely will
Let's find the time of flight
mv2/2=mgh => h=v2/2g
Also we can write that h=gt^2/2 whence t=sqrt(2h/g)
So the total time of flight T=2sqrt(2h/g)=2sqrt(2v^2/2g^2)=2v/g=2*20/10=4s